Opened 4 years ago

Closed 3 years ago

#588 closed flight processing (fixed)

EUFAR15/28, flight day 175a/2015, Mont Blanc

Reported by: stgo Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by stgo)

Data location: ~/arsf_data/2015/flight_data/france/EUFAR15_28-2015_175a_Mont_Blanc

Data arrived from ARSF via SATA disk on 09/07/2015

Scientific objective: Hyperspectral airborne measurements over snow, glaciers and vegetation; intercomparison with ground-based
measurements and modelling.

Priority: Unknown

PI: Antoine Rabatel

EUFAR Project ID: ISOTHERM

Sensors

  • Fenix (requested)
  • Leica LIDAR (requested)
  • Leica FW LIDAR (not requested, not flown)
  • Owl (not requested but flown anyway)

Change History (36)

comment:1 Changed 4 years ago by stgo

Navigation processing
Starting

comment:2 Changed 4 years ago by stgo

Navigation Processing

basestation ppp:

lat 45 44 21.34921
long 7 21 53.26501
height 598.474

comment:3 Changed 4 years ago by stgo

Navigation Processing
Good solution found with Ipas TC

comment:4 Changed 3 years ago by stgo

Fenix Processing

Lines 5, 6, 9, 11, 21 and 22 are all affected by identical GPS start/end points and times. The new end times have been calculated by dividing the lines by the sensor framerate and adding this to the end time. Using no nav so GPS points have been ignored.

comment:5 Changed 3 years ago by stgo

Fenix Processing

Found SCT values, these are drastically effected by using no nav files resulting in large offsets:

Flightline FENIX
1 1.80
2 3.18
3 2.09
4 3.60
5 9.98
6 44.93
7 2.28
8 3.96
9 8.32
10 1.04
11 2.55
12 4.00
13 2.75
14 2.71
15 8.34
16 3.54
17 3.90
18 3.22
19 2.94
20 2.06
21 57.03
22 48.81
23 2.58

comment:6 Changed 3 years ago by stgo

Lidar Processing

There are two separate sites in the dataset, one has a consistent pitch and roll error, the other varies wildly.

Site 1 consists of lines:
still finding pitch and roll values for these

150624_085753
150624_090334
150624_090912
150624_091446
150624_092003
150624_092504
150624_093027
150624_093602
150624_094159
150624_095317

site 2 consists of lines:

pitch -0.00246
roll -0.00081
heading 0

150624_100029
150624_100628
150624_101145
150624_101732
150624_102254
150624_102840
150624_103420
150624_104055
150624_104711
150624_105443
150624_110026
150624_110631
150624_111258
150624_111902
150624_112423

comment:7 Changed 3 years ago by stgo

Lidar processing

Pitch and roll found for lines:
150624_091446
150624_092003
150624_092504
150624_093027
150624_093602
150624_094159
150624_095317

pitch -0.00246
roll -0.00091
heading 0

The following lines have variable pitch and roll:
150624_085753
150624_090334
150624_090912

comment:8 Changed 3 years ago by stgo

Pitch and roll found:

150624_085753

pitch -0.00226
roll -0.00131
heading 0

150624_090334

pitch -0.00111
roll -0.00206
heading 0

The pitch and roll cannot be confirmed for 150624_090912 as it is not close enough to other flightlines to compare.

comment:9 Changed 3 years ago by stgo

Lidar Processing

No waveform files available, continuing as discrete laser delivery.

comment:10 Changed 3 years ago by stgo

  • Description modified (diff)

comment:11 Changed 3 years ago by stgo

Lidar processing

Created discrete lidar delivery, ready for delivery check.

comment:12 Changed 3 years ago by asm

Lidar Delivery Check
Started.

-Logsheet does not need the last two pages which are for Eagle and Hawk. Already solved and replaced.
-Pitch and roll looks perfectly fine.
-Some points in the ground has been incorrectly classified as noise for flightlines 11-25, specially when there is noise in the sky. They need to be manually re-checked but probably only flightline 25 worth redoing the classification.
-demcompare masked mean 1.57
-Everything else including projection, screenshots and dem files are fine.

comment:13 Changed 3 years ago by stgo

Lidar Delivery

Noise is being caused by trees, will check further then reclassify.

comment:14 Changed 3 years ago by stgo

Lidar Delivery

Finished reclassification, ready for re check

comment:15 Changed 3 years ago by asm

Lidar Delivery Check

-Classification looks better now.
-Those flightlines that have been re-classified need ascii files to be re-created since the number of noisy points have changed.

Last edited 3 years ago by asm (previous) (diff)

comment:16 Changed 3 years ago by stgo

Lidar Delivery

Recreated ascii files, ready for another check

comment:17 Changed 3 years ago by asm

Lidar Delivery Check
-Renamed las1.0 folder to las1.2
-Changes have been made and everything looks OK now.

Ready for delivery.

comment:18 Changed 3 years ago by asm

Lidar Delivery
Mail sent to PI and waiting in order to let him select delivery method.

comment:19 Changed 3 years ago by asm

Lidar Delivery
Lidar delivered via FTP. Notification has been sent. Finished.

comment:20 Changed 3 years ago by asm

Fenix DC

Started
-Removed fodis directory.
-Need to add logsheet.
-Need to add readme file.
-Found wrong SCT value for flightline 11 (and also double-check flightline 09).

comment:21 Changed 3 years ago by stgo

Fenix processing

This really should have been added to the ticket already... Sorry.

I have replaced the original header files with the updated wavescale and processed through. I created a new delivery last week. Will address asm comments in a moment

comment:22 Changed 3 years ago by stgo

Fenix processing

Looks like I had the decimal point off for line 11. SCT should have been 25.5!

comment:23 Changed 3 years ago by stgo

Fenix Processing

Updated line 11 with the corrected SCT. Also moved an updated readme in with replaced fenix mosaic.

Copied in the logsheet.

Ready for re-DC

comment:24 Changed 3 years ago by asm

Fenix DC

Had a quick look. SCT value for flightline 11 has been solved but strangely other flightlines has been also changed (check filestamps):
-Flightline 05 - With wrong SCT values
-Flightline 11 - Screenshot seems OK

Moreover:
-Flightline 09 has an strange bubble
-Flightline 02 has no usable data (could be removed from delivery or a note should be added on readme)
-Flightline 01 (a note should be added on readme about its data quality)

comment:25 Changed 3 years ago by stgo

I think flight 02 has usable data. I'm unwilling to remove it due to its area of coverage.

Can you clarify on what you mean by strange bubble at all?

I can't work out what has happened to the SCT values, I only processed line 11 so I don't see how line 5 could have changed. It seems the SCT values I wrote on this ticket are correct but the ones in the config/sync file are very wrong. Will update and reprocess.

comment:26 Changed 3 years ago by asm

Both stgo and I agree that flightline 02 could have some usable data and a note will be added to the readme and agree that flightline 09 seems to have a small error on the SCT value.

comment:27 Changed 3 years ago by stgo

Fenix delivery
I have fixed all the SCT errors in this delivery. The numbers entered in the table above are not correct (most missing one digit, seems I mucked up writing the values down), will post updated table in a minute. Also fixed a couple of additional file naming issues and the readme flightline names after creating a new delivery.

Ready for a recheck!

comment:28 Changed 3 years ago by asm

Fenix DC

Started
-SCT values look fine.
-Removed fodis directory.
-Removed additional pages from logsheet for eagle and hawk
-A note should be added to readme about flightline 02 data quality.

Needs the spectra to be checked before zipping files.

comment:29 Changed 3 years ago by asm

Fenix DC

Spectra looks fine. A note should be added to the ticket to remark that quite a few areas masked out due to overflows in the mapped data. Creating zipped files for delivery.

comment:30 Changed 3 years ago by stgo

Fenix DC

I've added the above comments to the Readme. Ready for check.

comment:31 Changed 3 years ago by asm

Fenix DC

All issues solved. Mapped files has been zipped. Project is ready to go.

comment:32 Changed 3 years ago by asm

Fenix delivery

Fenix delivered using FTP, slot 20. Notification has been sent to PI.

comment:33 Changed 3 years ago by dac

Archiving

Started archiving

comment:34 Changed 3 years ago by dac

Archiving

Owl renaming needs to be fixed before archiving raw data to NEODC.

comment:35 Changed 3 years ago by dac

Archiving

Owl renaming fixed by asm. Started upload to NEODC.

comment:36 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available from NEODC.

Note: See TracTickets for help on using tickets.